OpenStudy (anonymous):

Joyce combines 9 atoms of hydrogen and 3 atoms of nitrogen to make 3 molecules of ammonia. Write a direct variation equation that relates atoms of hydrogen, H, to atoms of nitrogen, N

OpenStudy (anonymous):

Formula, a = k*h*n

OpenStudy (anonymous):

3=k*9*3

OpenStudy (anonymous):

a=3 h=9 n=3

OpenStudy (anonymous):

find k

OpenStudy (anonymous):

k=(1/9)

OpenStudy (anonymous):

Yes. So the Equation would be: a = (1/9)*h*n

OpenStudy (anonymous):

omg thank u soo much!!1